### v1

# Partonic distribution functions and amplitudes using tensor network methods

Abstract
Calculations of the parton distribution function (PDF) and distribution amplitude (DA) are highly relevant to core experimental programs as they provide non-perturbative inputs to inclusive and exclusive processes, respectively. Direct computation of the PDFs and DAs remains challenging because they are non-perturbative quantities defined as light-cone correlators of quark and gluon fields, and are therefore inherently time-dependent. In this work, we propose a uniform quantum strategy based on tensor network simulation techniques to directly extract these hadronic quantities from first principles using the matrix product state of the hadrons in the same setup. We present exemplary numerical calculations with the Nambu-Jona-Lasinio model in 1+1 dimensions and compare with available exact diagonalization and quantum circuit simulation results. Using tensor networks, we evaluate the PDF and DA at various strong couplings in the large-qubit limit, which is consistent with expectations at perturbative and non-relativistic limits.
1 Introduction
The detailed understanding of the internal structure of nucleons in terms of fundamental quarks and gluons remains a cornerstone of research in high-energy nuclear and particle physics AbdulKhalek:2021gbh ; Accardi:2012qut . These structures are typically described using parton distribution functions (PDFs) and their extensions, such as transverse momentum-dependent parton distribution functions (TMDs), generalized parton distributions (GPDs), and parton distribution amplitudes (DAs). These functions are central to connecting theoretical predictions with experimental measurements in high-energy scattering processes. For example, PDFs play a critical role in predicting cross sections for processes such as single inclusive jet production in proton-proton collisions at colliders like the Large Hadron Collider (LHC). On the other hand, parton distribution amplitudes are key quantities in exclusive processes such as $\gamma^{*}\gamma\to\pi^{0}$.
Over the years, two principal approaches have been developed to determine non-perturbative parton distributions, such as PDFs and DAs, providing critical tools for exploring the internal structure of nucleons. The first approach is global QCD analysis, which exploits the QCD factorization framework Collins:1989gx . Within this framework, cross sections of physical processes are expressed as a convolution of perturbative short-distance coefficients and non-perturbative parton distributions. By analyzing diverse experimental cross-section data sets in conjunction with state-of-the-art perturbative calculations, PDFs or DAs can be extracted. This method has proven highly effective in incorporating experimental inputs to achieve precise determinations of parton distributions.
The second approach is based on lattice QCD, a method that allows for the first-principles computation of the QCD path integral in a discretized finite-volume Euclidean space-time. Direct lattice calculations of PDFs and DAs have traditionally been challenging because these quantities are defined as light-cone correlators, which are inherently time-dependent and thus difficult to compute in a Euclidean framework. However, significant progress has been made in recent years with the development of methods to compute quasi-PDFs Ji:2013dva ; Ji:2020ect and pseudo-PDFs Radyushkin:2017cyf , which are related to the standard light-cone PDFs through rigorous theoretical frameworks. These breakthroughs have enabled the indirect extraction of PDFs from lattice computations, representing a major advance in the field Lin:2017snn . See other approaches such as the hadronic tensors Liu:1993cv , axuiliary quarks Detmold:2005gg , current-current correlators Braun:2007wv ; Ma:2017pxb , operator product expansion Chambers:2017dov , and light-front Hamiltonian Brodsky:2006uqa ; 1stBLFQ , as well as Refs Ji:2020ect ; Cichy:2018mum ; Constantinou:2020pek for a review on the progress of recent developments.
Quantum computing technology has undergone rapid development and may provide an alternative path toward calculating the structure of hadrons from first principles Bauer:2022hpo ; DiMeglio:2023nsa . The exponential Hilbert space and quantum simulation algorithms provide a natural and efficient theoretical framework to simulate quantum field theory Jordan:2011ci ; Jordan:2014tma ; Jordan:2017lea . Studying real-time evolution. such as the case for the PDF and DA, is particularly suited to quantum simulation Lamm:2019uyc . The idea is to encode the fermionic and gluonic degrees of freedom into spin Hamiltonians and simulate the real-time matrix elements for evaluating these distribution functions and amplitudes directly on the quantum circuit. Recently, exploratory studies using quantum computing for parton structures are booming: from low-dimensional (1+1)-dimensional field theories Lamm:2019uyc ; Perez-Salinas:2020nem ; LiTianyin:2021kcs ; Li:2022lyt ; LiTianyin:2024nod ; Banuls:2024oxa ; Barata:2024bzk ; Grieninger:2024cdl ; Grieninger:2024axp , worldline formulation Mueller:2019qqj , spacetime Wilson loop Echevarria:2020wct , to effective light-front Hamiltonian models Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp , these studies already exhibit interesting physics and potential. However, most quantum simulation for parton structure is still restricted in performing systematic calculation and limited in extending to the continuum, largely due to today’s noisy intermeidate-scale quantum infrastructure Preskill:2018jim .
In this work, we aim to leverage tensor network (TN) methods in computing hadron structures. Traditionally used in condensed matter physics, TN have become an increasingly promising tool of its own to study hadronic physics and high-energy physics Silvi:2017srb ; Banuls:2018jag . Owing to its similarity with quantum computing, TN, despite being a classical computation approach, is used as a natural testbed for quantum simulation, especially for low dimensional and low entanglement problems. In particular, TN is an efficient simulation strategy to simulate quantum-many body systems by representing the quantum states locally in the Hilbert space and their dynamics can often be bounded by an area-law of entanglement entropy Orus:2013kga ; mps-reps-gs . Importantly, TN is free from sign problems and reliable in accessing regimes inaccessible to Monte Carlo methods Banuls:2018jag . Here, we aim to compute both the PDF and DA in the same setting of the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions using efficient TN simulation algorithms. Having direct access to hundreds of qubits with TN, we present numerical simulation at large-qubit limit that is representative of the continuum limit. Specifically, we compare the behavior of the PDF and DA at various mass values and strong coupling strengths to make qualitative comparison with available results in perturbative QCD and the non-relativistic limits.
We organize this paper as follows. In Sect. II, we introduce the PDF and DA both in general and also in the context of the 1+1 dimension NJL model. In addition, we detail its mapping to qubits and introduce suitable TN algorithms. In Sect. III, we present and discuss numerical results obtained using TN simulation. In particular, we study the numerical convergence, mass/strong coupling dependence, and comparison with perturbative calculations. In Sect. IV, we summarize the paper and discuss possible future directions.
2 Theoretical setup and computational methods
2.1 Parton distribution functions and amplitudes
Parton distribution functions (PDFs) are essential for describing inclusive processes at large momentum transfer in high-energy scattering, as dictated by the QCD factorization theorem Collins:1989gx . For example, the PDF for quarks within collinear factorization describes the probability for finding a quark carrying a longitudinal momentum fraction at some resolution scale. These functions are fundamentally defined through the matrix elements of quark field operators Collins:2011zzd , providing a direct connection between partonic structure and scattering observables. The PDF can be expressed by,
$f_{q/h}(x)=$ $\int_{-\infty}^{\infty}\frac{\mathrm{d}t}{4\pi}e^{-itx\vec{n}\cdot\vec{P}}\bra{h(\vec{P})}\bar{\psi}(t\vec{n})W(t\vec{n}\leftarrow\vec{0})\vec{n}\cdot\vec{\gamma}\psi(\vec{0})\ket{h(\vec{P})}\,,$ (1)
where $x$ denotes the longitudinal momentum fraction carried by the quark of the hadron $h(\vec{P})$ with momenta $\vec{P}$, $\psi$ represents the quark field operator, and $W(t\vec{n}\leftarrow\vec{0})$ is the Wilson line ensuring gauge invariance of the operator along the light-cone coordinate $\vec{n}$. In principle, one should expect the PDF satisfies,
$\sum_{q_{i}}\int_{0}^{1}\mathrm{d}x\,xf_{q_{i}/h}(x)=1\,,$ (2)
which is the momentum sum rule for all the partons.
On the other hand, the light-cone distribution amplitude (DA) governs exclusive processes at large momentum transfer Lepage:1980fj ; Efremov:1979qk . It is formulated from the light-like vacuum-to-meson matrix elements. For a hadron $h$, its DA is generally defined as the light-like separated, gauge-invariant vacuum-to-meson matrix element,
$\phi_{h}(x)=$ $\frac{1}{f_{h}}\int_{-\infty}^{\infty}\mathrm{d}t\,\,e^{-it(x-1)\vec{n}\cdot\vec{P}}\bra{\Omega}\bar{\psi}(t\vec{n})W(t\vec{n}\leftarrow\vec{0})\vec{n}\cdot\vec{\gamma}\psi(\vec{0})\ket{h(\vec{P})}\,,$ (3)
[…]
Since both $f_{q/h}(x)$ and $\phi_{h}(x)$ are boost invariant, the computation of the PDF and DA can be conveniently carried out in the hadron rest frame. In this frame, where $\vec{P}=0$, the above expressions can be simplified. Specifically, one replaces the $|h(\vec{P})\rangle$ with $|h(0)\rangle$, and the momenta $\vec{P}$ with the mass $M_{h}$ of the specified hadron $h$ in the PDF and DA. Still, computational difficulty remains. In terms of the Hamiltonian matrix, both the hadron state preparation and the real-time evolution of the Wilson line and field operators inherently require exact diagonalization. The complexity for diagonalizing an $n$-dimensional Hamiltonian (not necessarily sparse) essentially scales as $\mathcal{O}(n^{3})$ and therefore direct computation becomes a formidable if not an impossible task for a full size Hamiltonian that is relevant for comparison with experiments [^1] Though computationally expensive, classical Hamiltonian calculations for hadrons that are in good agreement with experimental data do exist using light-front Hamiltonian in the leading Fock sectors with sparse diagonalization Li:2017mlw ; Lan:2019vui ; Tang:2019gvn ; Shuo_Bc ; Qian:2020utg .. The goal of this work is to utilize tensor network methods as an alternative path for calculating the PDF and DA directly from first principles.
2.2 Nambu-Jona-Lasinio model in 1+1 dimensions
Simulating the 3+1 QCD Hamiltonian for hadronic structures is presently very challenging; therefore in this work we focus on the Nambu-Jona-Lasinio (NJL) model in 1+1 dimensions to demonstrate the computation of the PDF and DA in a common footing. The same procedure can also be applied to hadronic tensors. The NJL model is an effective model for low-energy two-flavored QCD Nambu:1961tp ; Nambu:1961fr and its simplified and renormalizable version, the Gross-Neveu (GN) model Gross1974 , is an ideal candidate for illustrative calculations including chiral symmetry breaking Shi:2015ufa , chiral condensate Huang:2001yw , phase transitions Thies:2019ejd ; Czajka:2021yll , and so forth. We start with the Lagrangian density of the GN model with zero chemical potential, given by,
$\mathcal{L}$ $=\bar{\psi}(i\not{\partial}-m)\psi+g(\bar{\psi}\psi)^{2}\,,$ (4)
where $m$ is the fermion mass and $g$ is the dimensionless coupling strength. Its Hamiltonian density is obtained via the Legendre transformation,
$\mathcal{H}$ $=\bar{\psi}(i\gamma_{1}{\partial_{1}}+m)\psi-g(\bar{\psi}\psi)^{2}\,.$ (5)
Here, the gamma matrices are chosen as $\gamma_{0}=\sigma^{z}$ and $\gamma_{1}=-i\sigma^{y}$, same as our previous work Czajka:2021yll . The fermion field $\psi(x)$ is discretized in position space and represented using staggered fermion fields Borsanyi:2010cj at each site $x_{n}=na$ on the lattice,
$\psi(x=x_{n})=\left(\begin{array}[]{cc}&\ \rho(x=x_{n})\\ &\ \eta(x=x_{n})\end{array}\right)=\frac{1}{\sqrt{a}}\left(\begin{array}[]{cc}&\ \chi_{2n}\\ &\ \chi_{2n+1}\end{array}\right)\,,$ (10)
where we use even $N$ for the total number of qubits and thus $\bar{N}=N/2$ for the total number of lattice sites. The lattice spacing is $\delta x=x_{n}-x_{n-1}=a$ where $n$ represents index of the lattice site. By using the forward-backward derivative convention Czajka:2021yll ; Chakrabarti:2003wi for the partial derivative of the fermion field, we extract the full Hamiltonian in staggered fermion representation,
$\begin{split}H=&-\frac{i}{2a}\left[\sum_{n=0}^{N-2}\left(\chi_{n}^{\dagger}\chi_{n+1}-\chi_{n+1}^{\dagger}\chi_{n}\right)\right]+m\sum_{n=0}^{N-1}(-1)^{n}\chi_{n}^{\dagger}\chi_{n}\\ &-\frac{g}{a}\sum_{n=0}^{N/2-1}\left(\chi_{2n}^{\dagger}\chi_{2n}-\chi_{2n+1}^{\dagger}\chi_{2n+1}\right)^{2}\,.\end{split}$ (11)
Here, we omit the periodic boundary terms $\chi_{N-1}^{\dagger}\chi_{0}-\chi_{0}^{\dagger}\chi_{N-1}$ in the kinetic term. In large systems, the long-range term has limited contribution and is complicated in the quantum simulation. To map the Hamiltonian onto quantum circuits, we apply the Jordan-Wigner (JW) transformation to the staggered fermion fields Jordan:1928wi ,
$\chi_{n}=\sigma_{n}^{-}\prod_{i=0}^{n-1}\left(-i\sigma_{i}^{z}\right),\quad\chi^{\dagger}_{n}=\sigma_{n}^{+}\prod_{i=0}^{n-1}\left(i\sigma_{i}^{z}\right)\,,$ (12)
and obtain the spin Hamiltonian,
$\begin{split}H&=\frac{1}{2a}\left[\sum_{n=0}^{N-2}\left(-\sigma^{+}_{n}\sigma^{z}_{n}\sigma^{-}_{n+1}+\sigma^{+}_{n+1}\sigma^{z}_{n}\sigma^{-}_{n}\right)\right]+m\sum_{n=0}^{N-1}(-1)^{n}\sigma^{+}_{n}\sigma^{-}_{n}\\ &-\frac{g}{a}\sum_{n=0}^{N/2-1}\left(\sigma^{+}_{2n}\sigma^{-}_{2n}-\sigma^{+}_{2n+1}\sigma^{-}_{2n+1}\right)^{2}\,,\end{split}$ (13)
where $\sigma_{i}^{x},\sigma_{i}^{y},\sigma_{i}^{z}$ are the Pauli matrices and $\sigma^{\pm}_{i}=(\sigma_{i}^{x}\pm i\sigma_{i}^{y})/2$ on qubit $i$.
Now we compute the PDF in eq. (1) and DA in eq. (3) using the obtained (1+1)-dimensional NJL Hamiltonian. Specifically, we use $\gamma^{+}=\frac{1}{\sqrt{2}}(\gamma^{0}+\gamma^{1})=\frac{1}{\sqrt{2}}(\sigma^{z}-i\sigma^{y})$ and $\vec{n}=(1,-1)$ for the light-cone coordinate. It follows that $\vec{n}\cdot{\vec{\gamma}}=\sqrt{2}\gamma^{+}$ and $\vec{n}t=(1,-1)z$ with a reparameterized variable $z$ for space and time. Since the gauge fields are absent in the NJL model, the explicit Wilson lines in the PDF and DA vanish and we are left with time evolution operators of the Hamiltonians. While this feature of the NJL model dramatically simplifies the calculation, we are well aware that one should include dynamical gluon contributions in a more realistic scenario with available methods for simulating the Wilson line Zohar:2012ts ; Brennen:2015pgn ; Lamm:2018siq ; Zohar:2019tsa ; Echevarria:2020wct , which we leave for a subsequent investigation. Applying the same JW encoding to field operators $\psi^{\dagger}$ and $\psi$, the final expression of the PDF and DA are related to the sum of four matrix element expectations,
$f_{q/h}(x)=$ $\frac{1}{4\pi}\sum_{z=\bar{N}_{\mathrm{min}}}^{\bar{N}_{\mathrm{max}}}e^{-ixM_{h}z}\bra{h}\big{(}M_{00}(z)+M_{11}(z)-M_{01}(z)-M_{10}(z)\big{)}\ket{h}\,,$ (14)
$\phi_{h}(x)=$ $\frac{1}{f_{h}}\sum_{z=\bar{N}_{\mathrm{min}}}^{\bar{N}_{\mathrm{max}}}e^{-i(x-1)M_{h}z}\bra{\Omega}\big{(}M_{00}(z)+M_{11}(z)-M_{01}(z)-M_{10}(z)\big{)}\ket{h}\,,$ (15)
where each matrix element operator is defined by,
$M_{ij}(z)$ $=e^{iHz}\chi^{\dagger}_{-2z+i+\bar{N}}e^{-iHz}\chi_{j+\bar{N}}\,,$ (16)
with evolution time $z=na/2$ relating to sites $n$ and $n+1$, and $\bar{N}_{\mathrm{min}}$ and $\bar{N}_{\mathrm{max}}$ mark the boundary of the lattice. Here, we always center the system around $x=0$ positions for the simulation, so $\bar{N}_{\mathrm{min}}=-\lfloor\bar{N}/2\rfloor$ and $\bar{N}_{\mathrm{max}}=-\lfloor\bar{N}/2\rfloor+\bar{N}-1$ with the center qubit index $\bar{N}$. Then, for a sufficiently large $\bar{N}$, our setup is a representation of the physical system. In practice, when calculating the PDF, we also subtract the vacuum expectation from the hadron expectation due to finite system size used in our simulation Collins:2011zzd .
2.3 Tensor network methods
Quantum computing technology emerges as one promising strategy to overcome classically difficult problems. Real-time evolution takes advantage of the trotter formula nielsen_chuang_2010 , in principle providing an efficient way to compute the PDF and DA. Various quantum algorithms for obtaining the ground state and excited states also provide solid approaches to state preparation wiersema2020exploring ; Choi:2020pdg ; Chakraborty:2020uhf ; Lin:2020zni ; Dong:2022mmq ; Kane:2023jdo . The final observables of interests are essentially expectation values that can be evaluated on quantum circuits via direct measurement and the Hadamard test nielsen_chuang_2010 . In this work, we focus on the calculation of the PDF and DA in the context of tensor network (TN) methods that share many similarities with quantum computation.
Tensor networks provide a convenient alternative for many-body simulations, especially since today’s quantum computers are still limited both in the number and fidelity of available qubits. TN methods recast (either exactly or up to accurate approximation) many-body quantum systems as a matrix product state (MPS) mps-reps-gs and use various algorithms and variational methods for studying the dynamics of MPS systems. MPS’s are especially well suited for 1D spatial systems as entropy area laws force an upper bound on entanglement vidal-erg . For this reason, TN methods provide efficient many body quantum simulations with classical hardware, especially for calculating hadronic structures in the 1+1 dimensional model considered in this work [^2] Note, for the same reason, TN do not replace standard QC methods but complement them..
2.3.1 Density matrix renormalization group
Preparation of the ground state or excited states is often the starting point for any quantum simulation Jordan:2011ci . While many quantum simulation methods have been suggested, including adiabatic quantum computation Chakraborty:2020uhf , Hamiltonian variational algorithm wiersema2020exploring , quantum subspace expansion mcclean2017hybrid , Rodeo algorithm Choi:2020pdg , among others; it is still typically very complicated to prepare the states on quantum circuits. In TN, by comparison, this complication is largely reduced utilizing efficient representation of the Hilbert space using MPS. An MPS is a many-body wave function decomposed into local tensors via singular value decomposition. In this work, we use the density matrix renormalization group (DMRG) algorithm to obtain the ground state and first excited state in MPS form White:1992zz ; White:1993zza ; Schollwock2005 ; schollwock2011density ; frank-dmrg .
The DMRG method basically takes the ansatz MPS and calculates the lattice configuration with the least amount of energy. It starts with an initial wave function, and then the energy is minimized by freely moving around two lattice sites while fixing all others. This process is repeated for many new lattice sites until convergence occurs. Traditionally, DMRG is used to study ground states of one-/two-dimensional many body systems frank-dmrg . The computational time scaling of the DMRG is $\mathcal{O}(N\chi^{3})$, where $\chi$ is bond dimension and $N$ is system size, i.e., total number of qubits dmrg-on-tpul . Specifically, the bond dimension is the dimension of the connecting index between two local tensors of an MPS. For calculating the first excited state, we construct a modified Hamiltonian that consists of the original Hamiltonian and a projection of the previously calculated ground state wave function. The projection has a coefficient “penalty” term that penalizes overlaps between the excited and ground states, which ensures orthogonality and minimization of the DMRG outside of the ground state ITensor:2022 . In this way, we use the DMRG algorithm to obtain both the ground state and the first excited state as MPS for the NJL model.
2.3.2 Time dependent variational principle
To calculate the matrix elements for the PDF and DA, as in eq. (16), we evolve the prepared MPS state in real time and collapse with a final MPS. The two primary algorithms for the time evolution of the MPS are the time-evolving block decimation (TEBD) Vidal:2003lvx ; verstraete2004matrix and time dependent variational principle (TDVP) Haegeman:2011zz ; haegeman2016unifying . Both of them have the same computational time scaling of $\mathcal{O}(N\chi^{3})$ as the DMRG algorithm, where $\chi$ is the bond dimension and $N$ is the system size. For time step $\tau$, the TDVP and TEBD algorithms have $O(\tau^{2})$ error scaling for the lower order dmrg-on-tpul ; tdvp-scaling ; vidal-tebd-many-body .
In principle, the TEBD method breaks up the Hamiltonian into small chunks and applies imaginary time evolution via Trotter gates that account for near neighbor interactions. The entanglement grows exponentially with time and TEBD method loses accuracy for long time evolution vidal-tebd-qc ; vidal-tebd-many-body . Specifically, for an $N$-qubit system, TEBD algorithm requires $\mathcal{O}(N\exp{(E_{\chi})})$ parameters where $E_{\chi}$ is the entanglement entropy, that is defined as von Neumann entropy $E_{\chi}=-\operatorname{Tr}(\rho\log(\rho))$ of the reduced density matrix $\rho$ from the whole quantum system to a block of local sites vidal-erg . If the bond dimension is not allowed to grow to account for this, then truncation errors occur vidal-tebd-qc ; vidal-tebd-many-body .
On the other hand, the TDVP algorithm projects the MPS onto a tangent space manifold, on which it solves the time dependent Schrodinger equation. This method doesn’t have the entanglement and error growth that Trotter gates have for larger systems, as it does not use Trotter gates at all. Instead, TDVP solves the Schrodinger equation projected onto a manifold, often chosen via the global Krylov method, which globally maintains all interaction terms Haegeman:2011zz . Potential errors are not as controllable, but are reduced with the use of Krylov time evolution itensor-tdvp . Though the TDVP and TEBD algorithms have the same scaling up to a prefactor, the TDVP algorithm has both a smaller prefactor for computational time scaling and less bond dimension growth Haegeman:2011zz ; itensor-tdvp . In this work, we primarily used TDVP, as it showed better performance than TEBD for NJL model, espeicially for high qubit simulations.
3 Results
In this work, we prepare the hadron state and simulate the real-time correlators of the quark and gluon fields using tensor network framework ITensor ITensor:2022 . The numerical results are compared with available exact diagonalization and quantum simulation results. Throughout this paper, we always use the lattice spacing $a=1$ in our simulations.
3.1 Hadron state preparation
Here, within the 1+1 dimension NJL model, our hadronic states are prepared via the DMRG algorithm. To prepare the hadron state on a tensor network within the neutral charge $Q=0$ sector, mimicking a charge-neutral pion ($\pi^{0}$), we utilize a charge-preserving (CP) lattice. With our setup, the total charge operator is defined as $Q=\sum_{i=1}^{N}(\sigma^{z}_{n}+(-1)^{n})/2$ on a lattice of $N$ qubits Ikeda:2024rzv . It is important to note that when working on a CP lattice, the Hamiltonian operator should be written in a form where its terms consist of only $\sigma^{z}$ or both $\sigma^{-},\sigma^{+}$ to ensure charge conservation, as in eq. (13). We first obtain the vacuum state (i.e. the ground state) on a CP lattice with sufficient number of sweeps and numerical accuracy. Specifically, we found 100 sweeps and 200 bond dimensions sufficient for simulating a system of 100 qubits. The high sweeps are found to be crucial to ensure DMRG convergence for simulation at larger qubits.
To obtain the hadron state (i.e., the first excited state), we project out the vacuum MPS and perform another set of DMRG sweeps with sufficiently large weights. For the hadron state, we found that even higher sweeps and maximum bond dimension are necessary to reach convergent mass results. Specifically, we used 500 sweeps and 800 bond dimensions. Throughout our simulation, we ensured that the obtained MPS consistently maintained a total charge of zero and that the bond dimension remained within the allocated maximum limit. In Table. 1, we summarize our hadron mass $M_{h}$, defined as the mass difference between the excited state and the ground state, findings for various system sizes $N$:
$M_{h}=\langle h|H|h\rangle-\langle\Omega|H|\Omega\rangle\,,$ (17)
Specifically, each value of $N$ is picked such that the two boundaries are matched up, i.e., $\bar{N}_{\mathrm{max}}=-\bar{N}_{\mathrm{min}}$. As the system size increases, $M_{h}$ becomes stable and the DRMG algorithm is able to identify the mass gap with high accuracy up to $10^{-8}$. For this reason, we always use $N=102$ qubits in preparing the ground and excited states for the final results.
Table 1: Hadron masses obtained from DMRG on CP lattice for different qubits. Here, $m=0.7a^{-1}$ and $g=0.25$. The DMRG results are accurate up to $10^{-8}$ with sufficient sweeps (100-500) and bond dimensions (200-800) at each qubit size.
Qubits (N) $18$ $30$ $50$ $102$
Hadron mass $M_{h}a$ 1.64600864 1.64359027 1.64246829 1.64187901
3.2 Real-time evolution
To compute the expectation of each matrix element as in eq. (16) in the PDF and DA, we apply the local operators, then real-time evolve the obtained MPS, and lastly collapse on the other MPS state to compute their overlap. Essentially, we repeatedly evaluate the following expectation,
[…]
where |$\ket{\psi_{L}},\ket{\psi_{R}}$ can be either |$\ket{h}$ or |$\ket{\Omega}$. Since |$\ket{\psi_{L}}$ is an eigenstate of $H$ (with eigenvalue $E_{L}$), the left-side time evolution operator reduces to an overall phase factor that greatly simplifies the computation [^3] We verified numerically the two expressions for $E_{ab}(z)$ in eq. (18) yield identical results.. Applying the local operators on MPS is straightforward and directly built into ITensor that involves three designated procedures: contraction, unpriming, and insertion of the MPS tensor. To time evolve the MPS, we use the TDVP algorithm Haegeman:2011zz ; haegeman2016unifying , included with the ITensorTDVP library. Specifically, we always use one time sweep and a maximum bond dimension of 810 (slightly larger than our max hadron state bond dimension). We tested the effect of different number of TDVP sweeps on the simulation results and found no significant improvement with more than one sweep. For large system sizes, we note that both the bond dimension of the MPS state and cost of the simulation increase at longer evolution time $z$. To avoid hitting the allocated maximum bond dimension and reduce computational complexity, we simulate in the region between $z=-10$ and $z=10$ where the matrix elements have substantial non-zero contributions. We verified that simulating beyond this region yielded no significant improvement to our final results.
Besides using the TDVP algorithm, we also implemented time evolution using the TEBD approach. At its heart, TEBD implements the Suzuki trotterization with approximation in the MPS. In comparison with the TDVP, we found it to be less accurate and more computationally demanding.
3.3 Parton distribution function and amplitudes
With our prepared hadron state and real-time evolution, we extract expectations of hadron matrix elements and then Fourier transform to obtain the PDF in eq. (14) and DA in eq. (15). Specifically, after obtaining the matrix element sums for various $z$ values, we interpolate them using cubic polynomial interpolator from scipy to generate a smooth curve prior to their Fourier transform (FT). For the FT, we used both the discrete and continuous FTs where we find no dramatic difference between them. For the numerical results presented below, we always show the continuous FT calculated using the composite Simpson’s rule from scipy. In addition, we use the momentum sum rule of the PDF and zero moment of the DA to normalize our obtained distributions after the FT respectively. In the case of the PDF, we also subtract the vacuum expectation of the matrix elements from the hadronic expectation before performing the FT.
For small system sizes, we find consistent agreement among the TN approach, exact diagonalization (ED), and quantum circuit (QC) simulation. For the quantum circuit simulation, we used first-order trotterization Trotter1959 to implement time evolution with scaling $O(\tau^{2})$ for the time step $\tau$ that matches the order of TEBD and TDVP childs2019nearly ; childs2021theory . State preparations for the QC simulation is dramatically simplified and taken as exact eigenstates. In Fig. 1, we present the numerical results for the PDF using a lattice of $N=10$ qubits. From Fig. 1(a), we see a strong alignment among all three results, implying that our setup is valid and generally applicable to quantum computing. In Fig. 1(b), we make a log-linear plot of the absolute error $\Delta f_{q}(x)$ in TN and QC methods compared with exact diagonalization for each $x$. Noticeably, the TN consistently and significantly reduces the error by six orders of magnitude with an average error of around $10^{-8}$, which aligns with the numerical accuracy of the DMRG method for state preparation. Similar observations are found in computing the DA at $N=10$ qubits. These observations prove the reliability of the TN strategies used in our numerical calculations.
(a)  PDF calculated using different methods
(b)  Errors in PDF compared to ED
Figure 1: Simulation results for the PDF comparing tensor network, exact diagonalization, and quantum computing methods. We use $N=10$, $m=0.7a^{-1}$, and $g=0.25$. Right plot shows the error of the PDF calculated using tensor network and quantum circuit simulation.
To demonstrate numerical results in the continuum limit, we use sufficiently high number of qubits in our simulations to compute the PDF and DA using tensor networks. Such calculation is beyond the current limits of exact simulation and quantum simulation with classical hardware. With tensor networks, we are able to study physical results at continuum limits utilizing an efficient representation of the many-body system with the MPS. In Fig. 2, we compute both the PDF and DA in the same setting with increasing system sizes (i.e., qubits) using selected mass $m=0.7a^{-1}$ and strong coupling $g=0.25$. For the PDFs, as expected in a one-flavor hadron (such as $\pi^{0}$, composed of many pairs of $q$ and $\bar{q}$), we see an asymmetric plot with a peak at $x=0.5$ and $x=-0.5$. Similarly for the DAs, we see a typical distribution centered in the $x\in[0,1]$ region. As system size $N$ increases, both the PDF and DA results gradually converge to the large-qubit curve at $N=102$, suggesting the calculations approach the continuum of $N\rightarrow\infty$ limit [^4] In principle, one would take both $a\rightarrow 0$ and $N\rightarrow\infty$. For simplicity, we fix $a=1$.. Therefore, for all the results below, we always use $N=102$ qubits for the simulation and we find it sufficient to represent the continuum behavior of the PDF and DA at this large qubit limit for our NJL Hamiltonian. Overall, the shape and tendency of our results are qualitatively consistent with experimental observations such as E791 data E791:2000xcx for the PDF and Fermilab E615 data Conway:1989fs for the DA of the pion.
(a)  PDF
(b)  DA
Figure 2: Simulation results for the PDF and DA over increasing number of qubits $N$. Here, we fix $m=0.7a^{-1}$ and $g=0.25$. The asymptotic of perturbative QCD for the DA is provided.
One important strength of our approach is the freedom to investigate the PDF and DA at various coupling strengths, especially in the non-perturbative regions that are generally complicated. In Fig. 3, we show the PDF and DA results at various coupling strengths $g=0.25,0.5,1.0$ with the same $m=0.5a^{-1}$ and $N=102$ qubits. Interestingly, the PDFs show distinctive behavior at different coupling strengths. At small coupling $g=0.25$, we get a sharp peak at $x=0.5$ and $x=-0.5$ that is consistent with the perturbative QCD expectation. With increased coupling strength, the PDF becomes less sharply peaked. Similarly for the DA, at the low coupling such as $g=0.25$, the distribution is sharply centered at $x=0.5$, comparable to the perturbative limit of $\phi_{\mathrm{asym}}(x)=6x(1-x)$ Lepage:1980fj . When we turn on the coupling strength towards $1$, we see significant non-perturbative effects, giving broadened DA curves. Qualitatively, our results are also in agreement with available quantum and tensor network simulation results from the light-front Hamiltonian methods Kreshchuk:2020dla ; Kreshchuk:2020aiq ; Qian:2021jxp and lattice calculations LiTianyin:2021kcs ; LiTianyin:2024nod ; Banuls:2024oxa .
(a)  PDF
(b)  DA
Figure 3: Simulation results for the PDF and DA at different values of strong coupling $g$. The mass is fixed at $m=0.5a^{-1}$ and total qubits are $N=102$.
Likewise, we compute the PDF and DA for various mass values which may serve to illustrate the separated behavior between heavy systems and light systems. In Fig. 4, we compute the PDF and DA at masses $m=0.1a^{-1},0.25a^{-1},0.5a^{-1}$ with fixed $g=0.25$ and $N=102$ qubits. In both the PDF and DA, we find a more peaked distribution at $x=0.5$ for the heavier masses and broadened distributions for the lighter masses. We also observe less oscillations on the tail ends of the distributions as mass decreases. Importantly, our findings are comparable to the light-front Hamiltonian approach Brodsky:1997de ; Brodsky:2014yha ; 1stBLFQ for mesonic systems of distinctive masses, from light mesons Vega:2009zb ; Swarnkar:2015osa ; Lan:2019vui ; Jia:2018ary ; Qian:2020utg to heavy quarkonia Swarnkar:2015osa ; Li:2017mlw ; Lan:2019img , where leading Fock sectors are used. It would be interesting to further address the large-qubit calculations for these different physical systems by including flavor degrees of freedom into our Hamiltonian.
(a)  PDF
(b)  DA
Figure 4: Simulation results for the PDF and DA at different values of mass $m$. The strong coupling is fixed at $g=0.25$ and total qubits are $N=102$.
4 Conclusion and outlook
In this work, we proposed a quantum strategy based on tensor network simulation to extract the PDF and DA for hadrons in the Nambu-Jona-Lasinio model in 1+1 dimensions. Our approach utilizes the DMRG algorithm to accurately construct the hadron and vacuum states on a charge-preserving lattice and the TDVP algorithm to efficiently time evolve the two-point correlator needed for these non-perturbative quantities. With our approach, we extract the PDF and DA in the same setting with converging results up to 102 qubits. We are able to study the behavior of the PDF and DA under the same setting for various masses and strong couplings, and observe consistent agreement with the expectations from the perturbative QCD as well as non-relativistic limit. Our exploratory study shows the advantage of tensor network in the current era for computing various physically important quantities in the continuum limit.
It would be very important to expand non-perturbative calculations of quantities such as the PDF, DA, and hadronic tensors using tensor network to U(1) and SU(2) gauge theories with consideration of multiple spins and flavors. Inclusion of flavor degrees of freedom in the line of study would provide significant understanding to realistic mesons and baryons from first principles. While tensor networks have primarily been applied to 1+1 dimensions, it would be interesting to extend to higher lattice dimensions. The natural extension of MPS to higher dimensions would be projected entanglement pair states Verstraete:2004cf ; verstraete2004valence , tree tensor network Shi:2005ctp , and so forth, which have been used to study 3+1 dimension Lattice Gauge Theory Magnifico:2020bqt ; Banuls:2018jag . We also plan to incorporate more advanced simulation techniques, such as tensor renormalization group LevinNave:2007 and its natural extension, tensor network renormalization tnr , to scale to higher qubit simulations with reduced error and study entanglement structure. These methods can be used to calculate correlation functions and generalize to systems beyond one spatial dimension.
Acknowledgement
We are grateful to João Barata, Kazuki Ikeda, Tianyin Li, Yang Li, David Lin, Manuel Schneider, and Hongxi Xing for their helpful and valuable discussions. ZK, NM, and PN are supported by National Science Foundation under grant No. PHY-1945471. WQ is supported by the European Research Council under project ERC-2018-ADG-835105 YoctoLHC; by Maria de Maeztu excellence unit grant CEX2023-001318-M and project PID2020-119632GB-I00 funded by MICIU/AEI/10.13039/501100011033; by ERDF/EU; by the Marie Sklodowska-Curie Actions Fellowships under Grant No. 101109293; and by Xunta de Galicia (CIGUS Network of Research Centres).